# Torque-based Graph Surgery: Enhancing Graph Neural Networks with Hierarchical Rewiring

### v1

@@ 6 Conclusion
In summary, we proposed a Torque-driven hierarchical rewiring graph neural network (TorqueGNN), which dynamically refined the graph structures to enhance representation learning on heterophilous and noisy graphs. By introducing an interference-aware torque metric, the product of the distance and the energy between node representations, TorqueGNN automatically removed undesirable connections and introduced beneficial ones during message passing. This hierarchical rewiring yielded interference-resilient, importance-aware propagation tailored to each layer’s receptive field. Extensive evaluations across homophilous and heterophilous benchmark datasets demonstrated that TorqueGNN consistently outperformed state-of-the-art baselines and maintained robust performance on adversarial graphs. We state the limitations of our model in Appendix D.
Appendix A Algorithm
Algorithm 1 outlines the complete workflow of TorqueGNN.
Input: Node features $\{\mathbf{x}_{i}\in\mathbb{R}^{d}\}_{i=1}^{N}$, candidate edge set $\mathcal{T}$, ground truth matrix $\mathbf{Y}$, hyperparameters $L$ and $\alpha$.
@@ 2 $\mathbf{h}_{i}^{(0)}=\text{ReLU}(\mathbf{x}_{i}\mathbf{\Theta})$;
5   Compute pairwise distance $D_{\langle i,j\rangle}^{(l)}$ and energy $E_{\langle i,j\rangle}^{(l)}$ with Eqs. (6) and (7);
6   Compute the $l$th order torques with Eq. (5) and sort them. // Torque computation
7   Gain the largest torque gap $\mathcal{K}$ with Eqs. (8)-(9);
@@
9   Compute the sampling probability of candidate edges with Eq. (11);
@@
11   Update node representation $\mathbf{h}_{i}^{(l)}$ with Eq. (12). // Message passing
@@
14 for $t=1\rightarrow T$ do
[…]
16
17Obtain $\mathbf{\hat{y}}_{i}=\text{Softmax}\left(\mathbf{h}^{(L)}_{i}\mathbf{\Phi}\right)$;
@@
Algorithm 1 TorqueGNN
Appendix B Complexity Analysis
The dominant computational cost of TorqueGNN lies in: 1) Torque computation and graph rewiring. For each order $l$, we compute torque values only on the edges in $\mathcal{A}^{(l)}$, costing $\mathcal{O}(|\mathcal{A}^{(l)}|)$, and then sort these values in $\mathcal{O}(|\mathcal{A}^{(l)}|\log|\mathcal{A}^{(l)}|)$. When adding edges, if the candidate set size is $B$, the combined probability calculation and sorting cost is $\mathcal{O}(B+B\log B)$. 2) Message passing on the rewired graph $\mathcal{A}^{(l)}$. Initial representations $\mathbf{H}^{(0)}$ are obtained by a fully connected layer with parameter $\mathbf{\Theta}\in\mathbb{R}^{d\times m}$ on $\mathbf{X}\in\mathbb{R}^{N\times d}$, at cost $\mathcal{O}(Ndm)$. Aggregation over $\mathcal{A}^{(l)}$ then costs $\mathcal{O}(m|\mathcal{A}^{(l)}|)$ per layer. The output layer with $\mathbf{\Phi}\in\mathbb{R}^{m\times c}$ requires $\mathcal{O}(Nmc)$. Putting these together for an $L$-layer network and assuming $B\ll|\mathbf{A}^{l}|$ for all $l$, the overall complexity is $\mathcal{O}(LNdm+\sum_{l=1}^{L}|\mathcal{A}^{(l)}|\log|\mathcal{A}^{(l)}|)$.
@@ C.2 Datasets
    • Citeseer comprises 3,327 publications classified into six categories, with each paper encoded by a 3,703-dimensional binary word-presence vector.
@@
C.3 Methods
@@ C.3.1 GNNs for Homophilous and Heterophilous Graphs
GAT leverages attention mechanisms to weight the influence of neighboring nodes during message passing. Its core rule is:
$\mathbf{h}_{i}^{(l+1)}=\sigma\left(\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}\mathbf{W}\mathbf{h}_{j}^{(l)}\right),\alpha_{ij}=\frac{\exp\left(\operatorname{LeakyReLU}\left(a^{\top}\left[\mathbf{W}\mathbf{h}_{i}^{(l)}\|\mathbf{W}\mathbf{h}_{j}^{(l)}\right]\right)\right)}{\sum_{k\in\mathcal{N}_{i}}\exp\left(\operatorname{LeakyReLU}\left(a^{\top}\left[\mathbf{W}\mathbf{h}_{i}^{(l)}\|\mathbf{W}\mathbf{h}_{k}^{(l)}\right]\right)\right)}.$ (20)
@@
H2GCN contains three designs, and the first design aims to separate ego- and neighbor-message:
$\mathbf{h}_{i}^{(l)}=[\mathbf{h}_{i}^{(l-1)}||\text{Agg}(\{\mathbf{h}_{j}^{(l-1)}:v_{j}\in\mathcal{N}_{i}\})]$ (23)
Second, H2GCN directly aggregates neighbors of different hops, which enhances high-order information:
$\mathbf{h}_{i}^{(l)}=[\mathbf{h}_{i}^{(l-1)}||\text{Agg}(\{\mathbf{h}_{j}^{(l-1)}:v_{j}\in\mathcal{N}^{1}_{i}||\text{Agg}(\{\mathbf{h}_{j}^{(l-1)}:v_{j}\in\mathcal{N}^{2}_{i}||\cdots\})].$ (24)
Finally, the output is gained by combining the intermediate representations of all layers:
$\mathbf{h}_{i}=[\mathbf{h}_{i}^{(1)}||\mathbf{h}_{i}^{(2)}||\cdots||\mathbf{h}_{i}^{(L)}].$ (25)
@@
FAGCN adaptively balances low- and high-frequency graph signals via a learnable mixing coefficient:
$\mathbf{h}_{i}^{(l)}=\eta\mathbf{h}_{i}^{(0)}+\sum_{j\in\mathcal{N}_{i}}\frac{\alpha_{ij}}{\sqrt{d_{i}d_{j}}}\eta\mathbf{h}_{i}^{(l-1)},$ (27)
where $\alpha_{ij}\in(-1,1)$ is a learnable coefficient.
ACM-GNN dynamically learns how to weight multiple propagation channels:
$\mathbf{H}^{(l)}=\operatorname{ReLU}\left(\operatorname{diag}\left(\alpha_{H}^{(l)}\right)(\mathbf{I}-\mathbf{P})\mathbf{H}^{(l-1)}\mathbf{W}_{H}^{(l)}+\operatorname{diag}\left(\alpha_{L}^{(l)}\right)\mathbf{P}\mathbf{H}^{(l-1)}\mathbf{W}_{L}^{(l)}+\operatorname{diag}\left(\alpha_{I}^{(l)}\right)\mathbf{H}^{(l-1)}\mathbf{W}_{I}^{(l)}\right).$ (28)
Here, $\alpha$ is adaptive mixing weight.
OGNN introduces an order-aware message passing by sorting neighbors according to learned importance scores before aggregation:
$\mathbf{h}_{i}=\mathbf{h}_{i}^{(1)}\odot g_{i}^{(1)}+\mathbf{h}_{i}^{(2)}\odot\left(g_{i}^{(2)}-g_{i}^{(1)}\right)\ldots+\mathbf{h}_{i}^{(L)}\odot\left(g_{i}^{(KL}-g_{i}^{(L-1)}\right),$ (29)
where $g_{i}^{(l)}$ is a vector with a split point $p_{i}^{(l)}$ so that its first entries are 0 and its remaining entries are 1.
M2MGNN partitions neighbor embeddings into class-aligned “chunks” via learnable soft labels:
$\mathbf{h}_{i}^{(l)}=\operatorname{ReLU}\left((1-\beta)\mathbf{h}_{i}^{(0)}+\beta\mathbf{m}_{i}^{(l)}\right),\mathbf{C}_{it}^{(l)}=\sum_{j\in\mathcal{N}_{i}}\mathbf{s}_{t}^{(l)}(i,j)\hat{\mathbf{h}}_{j}^{(k)},\quad\mathbf{m}_{i}^{(l)}=\|_{t=1}^{C}\mathbf{C}_{it}^{(l)},$ (30)
where $\mathbf{s}_{t}^{(l)}(i,j)$ is the score of node $v_{i}$ towards $v_{j}$ in the $t$-th chunk.
PCNet introduces a unified two-fold filtering framework that simultaneously performs homophilous low-pass smoothing and heterophilous high-pass aggregation:
$\mathbf{H}=(\sum_{k=1}^{K}\theta_{k}\mathrm{PC}_{k}(\tilde{L})+\theta_{\mathrm{0}})\mathbf{X}\mathbf{W}.$ (31)
Here, $\mathrm{PC}_{k}(\tilde{L})$ denotes the $n$-th Poisson–Charlier polynomial of the normalized Laplacian $\tilde{L}$, and $\{\theta_{k}\}_{k=0}^{K}$ are learnable filter coefficients.
@@ C.3.2 Rewiring Strategies
FoSR develope a spectral-expansion–based rewiring algorithm that added edges to alleviate over-squashing, and integrate it with a relational GNN to provably avoid oversmoothing.
JDR simultaneously refine graph topology and node features by aligning the principal eigenspaces of their matrices, boosting GNNs across both homophilous and heterophilous networks.
DropEdge randomly remove edges at each training epoch to act as both data augmentation and message-passing reduction, which is used to mitigate over-fitting and over-smooting problems.
@@ C.4 Hyperparameters
In the subsection, we list the detailed hyperparameters used for the experiments and they are also provided in code. The hyperparameters can be found in Table 4.
Table 4: Hyperparameters of TorqueGNN in 10 datasets.
Datasets Lr Wd Dropout $L$ $t$ $r$ epochs
Texas 0.001 5e-4 0.7 8 5 0.2 100
Wisconsin 0.001 0.5 0.5 8 5 0.2 100
Cornell 0.001 0.05 0.7 8 2 0.2 100
Actor 0.001 0.05 0.1 8 5 0.2 100
Citeseer 0.001 5e-8 0.1 8 2 0.2 100
Cora 0.001 0.1 0.1 8 2 0.2 100
Pubmed 0.001 0.08 0.1 12 5 0.2 100
Tolokers 0.001 5e-8 0.1 2 2 0.2 500
Questions 0.001 5e-8 0.1 2 2 0.2 500
Penn94 0.001 5e-8 0.1 2 2 0.2 500
@@ C.5 Experiments
Ablation Study. To validate the effectiveness of our proposed torque metric, we conduct ablation studies comparing three edge-scoring schemes. As reported in Table 5, reliance on structural distance alone yields the poorest classification performance on most benchmarks, reflecting the fact that certain long-range connections convey meaningful semantics. Introducing the energy metric markedly improves accuracy, confirming its significance as a measure of edge perturbation strength. Moreover, on Tolokers and Pubmed, the energy-only variant achieves the highest accuracy, slightly surpassing the combined torque score, likely because some low-interference connections carrying critical semantic information are inadvertently pruned when the distance is too large. Crucially, our torque score, integrating both distance and energy, consistently outperforms either criterion besides Tolokers and Pubmed, thereby validating its superior ability to accurately identify disruptive edges, guide interference-aware rewiring, and enhance overall classification accuracy.
Table 5: Ablation Study: Performance of TorqueGNN equipped with different edge-scoring metrics; the best results are shown in bold.
Datasets Distance Energy Torque (Distance $\times$ Energy)
Texas 88.61 (0.70) 89.16 (0.69) 91.11 (1.67)
Wisconsin 79.20 (2.40) 84.60 (1.80) 89.60 (1.20)
Cornell 81.94 (4.17) 83.05 (0.83) 89.44 (1.67)
Actor 36.69 (0.40) 35.48 (0.07) 36.72 (0.16)
Tolokers 72.46 (0.04) 76.11 (0.28) 75.28 (0.02)
Penn94 81.11 (0.09) 81.20 (0.15) 81.30 (0.19)
Questions 63.11 (1.26) 62.65 (0.53) 63.54 (0.85)
Citeseer 73.35 (0.24) 76.53 (0.42) 76.84 (0.60)
Cora 78.30 (0.59) 84.50 (0.00) 85.39 (0.30)
Citeseer 73.35 (0.24) 76.53 (0.42) 76.84 (0.60)
Pubmed 88.01 (0.11) 88.52 (0.16) 88.36 (0.08)
Parameter Sensitivity. Figure 5 explores the effect of network depth $L$. Across most datasets, performance peaks at $L=8$, thereby validating the effectiveness of our THR strategy for mining high-order information. Beyond eight layers, accuracy declines, primarily due to training challenges typical of deep networks, such as over-smoothing. While our rewiring mechanism mitigates over-smoothing to some extent, it is not explicitly engineered to resolve this phenomenon; overcoming depth-related bottlenecks remains a direction for future work.
Figure 5: Parameter sensitivity: Performance curves on five datasets with layers changing in $2\}$ , 4 , 8 , 16 , 32 }.
Appendix D Broader Impact Statement and Limitations
@@
THR depends on sorting operations that impose considerable computational and memory burdens on large, densely connected graphs. Additionally, by assigning equal weight to distance and energy metrics, THR may inadvertently prune edges that carry essential semantic information and exhibit low interference when the distance is too large. How to optimally balance these two metrics in accordance with the characteristics of different datasets remains an open challenge.

### v3

@@ 7 Conclusion
In summary, we proposed a Torque‑driven hierarchical rewiring strategy (THR), which dynamically refined the graph structures to enhance representation learning on heterophilous and homophilous graphs. By introducing an interference‑aware torque metric, the product of the displacement vector and the feature vector weighted by the homophily ratio disparity, THR automatically removed undesirable connections and introduced beneficial ones during message passing. This hierarchical rewiring yielded interference‑resilient, importance‑aware propagation tailored to each layer’s receptive field. Extensive evaluations across homophilous and heterophilous benchmark datasets demonstrated that THR consistently obtained the performance gains and outperformed other rewiring methods.
Appendix A Appendix
Appendix B Algorithm
Algorithm 1 outlines the complete workflow of APPNP with THR.
Input: Node features $\{\mathbf{x}_{i}\in\mathbb{R}^{d}\}_{i=1}^{N}$, candidate edge set $\mathcal{T}$, ground truth matrix $\mathbf{Y}$, the number of layers $L$, hyperparameters $t$ and $\alpha$.
@@ 2 $\mathbf{h}_{i}^{(0)}=\text{ReLU}(\mathbf{x}_{i}\mathbf{\Theta})$;
5   Compute pairwise distance $D_{\langle i,j\rangle}^{(l)}$ and homophily disparity $E_{\langle i,j\rangle}$ with Eqs. 4 and 5;
6   Compute the $l$th order torques with Eq. 6 and sort them. // Torque computation
7   Gain the largest torque gap $\mathcal{K}$ with Eqs. 7-8;
@@
9   Compute the sampling probability of candidate edges with Eq. 10;
@@
11   Update node representation $\mathbf{h}_{i}^{(l)}$ with Eq. 11. // Message passing
@@
14
15Obtain $\mathbf{\hat{y}}_{i}=\text{Softmax}\left(\mathbf{h}^{(L)}_{i}\mathbf{\Phi}\right)$;
@@
Algorithm 1 GNN with THR
@@ C.2 Datasets
    • Citeseer comprises 3,327 publications classified into six categories, with each paper encoded by a 3,703-dimensional binary word‐presence vector.
@@
    • Flickr [55] is an undirected graph originated from NUS-wide, including 89,250 nodes and 2,724,458 edges. Each node is an image with 500-dimensional bag-of-word features and each edge links two images sharing some common properties.
C.3 Baselines
@@ C.3.2 Rewiring Strategies
DropEdge randomly remove edges at each training epoch to act as both data augmentation and message‑passing reduction, which is used to mitigate over-fitting and over-smooting problems.
FoSR is a preprocessing method, which aim to address the oversquashing issue by improving the graph connectivity. It adds edges by exploring the first order change in the spectral gap.
BORF uses the Ollivier-Ricci curvature to rewire graph, where minimally curved edges causing the information bottlenecks should add connections and maximally curved edges leading to over-smoothing should be removed.
SJLR combines the Jost–Liu Curvature of each edge with the embedding similarity between its incident nodes, and uses the weighted score as the probability for edge removal or addition.
DHGR compares the neighborhood feature distribution and neighborhood label distribution between node pairs; edges connecting nodes with low similarity (heterophilous) are pruned, while edges between highly similar (homophilous) nodes are added.
DHGR vs. THR. Although both methods essentially assess edge homophily or heterophily through feature and label differences, they follow distinct methodological lines. DHGR is a preprocessing approach that aggregates neighborhood features and derives local label distributions from pseudo-labels produced by a pre-trained model, a heuristic design without explicit theoretical grounding. By contrast, THR operates within the optimizing model, contrasting node representations and quantifying their homophily ratio disparity, thereby aligning with prior theoretical proofs and offering a more principled formulation.
@@ C.4 Hyperparameters
In the subsection, we list the detailed hyperparameters used for the experiments and they are also provided in code. The hyperparameters can be found in Tables 6-7.
Table 5: Hyperparameters of THR on GCN across 11 datasets.
Datasets Lr Wd Dropout L t epochs Normalize Data Hidden Size
Texas 0.05 0.0005 0.5 2 5 10000 Yes 32
Wisconsin 0.05 0.0005 0.5 2 5 10000 Yes 32
Cornell 0.05 0.0005 0.5 2 5 10000 Yes 32
Actor 0.01 0.0005 0.5 2 2 10000 No 32
Citeseer 0.01 0.0005 0.5 2 10 10000 Yes 32
Cora 0.01 0.05 0.5 2 2 10000 No 32
Pubmed 0.01 0.0005 0.5 2 2 10000 Yes 32
Tolokers 0.005 5e-8 0.2 2 1 10000 No 32
Questions 0.005 5e-8 0.2 5 1 10000 No 32
Penn94 0.001 5e-8 0.5 2 1 10000 No 32
Flickr 0.01 0.0005 0.5 2 1 10000 No 32
Table 6: Hyperparameters of THR on GPRGNN across 11 datasets.
Datasets Lr Wd Dropout L t epochs Normalize Data PPR Hidden Size
Texas 0.05 0.0005 0.5 2 5 10000 Yes 1 32
Wisconsin 0.05 0.0005 0.5 2 5 10000 Yes 1 32
Cornell 0.05 0.0005 0.5 2 5 10000 Yes 0.9 32
Actor 0.01 5e-8 0.5 2 2 10000 Yes 0.9 32
Citeseer 0.01 0.0005 0.5 2 10 10000 Yes 0.1 32
Cora 0.01 0.0005 0.5 2 5 10000 Yes 0.1 32
Pubmed 0.05 0.0005 0.5 2 2 10000 Yes 0.2 32
Tolokers 0.005 5e-8 0.5 2 1 10000 No 0.1 256
Questions 0.05 5e-8 0.5 2 1 10000 No 0.1 32
Penn94 0.01 0.0001 0.5 2 1 10000 No 0.1 32
Flickr 0.01 0.0005 0.5 2 1 10000 No 0.1 32
Table 7: Hyperparameters of THR on APPNP across 11 datasets.
Datasets Lr Wd Dropout $\alpha$ $L$ $t$ epochs Normalize Data Hidden Size
Texas 0.001 0.0005 0.7 0.05 8 5 100 No 512
Wisconsin 0.001 0.5 0.5 0.05 4 5 100 No 512
Cornell 0.001 0.05 0.7 0.5 8 2 100 No 512
Actor 0.001 0.05 0.1 0.5 8 5 100 No 512
Citeseer 0.001 0.05 0.4 0.5 8 10 100 No 512
Cora 0.001 0.5 0.4 0.5 4 2 100 No 512
Pubmed 0.001 5e-8 0.4 0.5 4 2 100 No 512
Tolokers 0.001 5e-8 0.1 0.8 2 2 500 No 512
Questions 0.001 5e-8 0.1 0.8 2 2 500 No 512
Penn94 0.001 5e-8 0.1 0.8 2 2 500 No 512
Flickr 0.001 0.5 0.1 0.8 2 2 500 No 512
@@ C.5 Experiments
Classification Results. Table 8 shows the performance gains brought by APPNP with diverse rewiring methods. We can observe that on most datasets, THR obtains the optimal performance, indicating its effectiveness.
Table 8: Node classification results on benchmark datasets with APPNP as the backbone models: Mean ACC % (Standard Deviation %). The first- and second-best accuracies are highlighted in bold and underlined, respectively.
Methods/Datasets APPNP FoSR BROF SJLR DHGR DropEdge-L THR
Texas 49.17±3.30 78.04 (3.70) 73.53 (7.66) 81.57 (3.94) 78.04 (4.62) 72.75 (4.68) 84.12 (3.22)
Wisconsin 47.60±4.54 74.26 (4.47) 75.00 (4.41) 83.53 (5.95) 73.82 (3.77) 71.91 (4.85) 87.79 (3.54)
Actor 35.24 (0.56) 35.51 (1.42) 35.35 (1.28) 35.19 (1.13) 35.90 (1.16) 35.48 (1.12) 36.34 (0.87)
Cornell 67.57 (5.54) 67.57 (5.54) 68.38 (7.46) 74.59 (5.16) 69.46 (8.80) 69.19 (7.27) 77.57 (8.02)
Penn94 76.53 (0.28) 76.53 (0.28) OoM 79.73 (0.24) 79.10 (0.40) 76.13 (0.40) 82.56 (0.43)
Flickr 57.26 (7.69) 56.31 (6.99) OoM 61.86 (5.17) 62.20 (1.10) 61.25 (3.30) 63.28 (1.56)
Citeseer 74.02 (0.38) 77.65 (1.55) 77.65 (1.24) 77.25 (1.35) 76.89 (1.81) 77.69 (1.67) 78.74 (1.29)
Cora 85.89 (1.19) 85.89 (1.19) 85.19 (1.87) 86.51 (1.59) 85.85 (1.79) 85.54 (1.14) 86.35 (1.61)
Pubmed 87.19 (0.55) 87.19 (0.55) 87.16 (0.40) 88.84 (0.40) 88.47 (0.44) 87.66 (0.33) 88.31 (0.49)
Tolokers 75.11 (0.74) 75.11 (0.74) OoM 78.46 (1.11) 75.33 (0.83) 74.64 (1.06) 79.29 (0.42)
Parameter Sensitivity. Although $\alpha$ balancing the contribution of the learned high-order representation and the original input features originates from APPNP, THR modifies the graph structure over which propagation occurs. To examine how signal diffusion changes with respect to $\alpha$ under the rewired graph, we perform a sensitivity analysis shown in Figure 4, where a larger $\alpha$ increases the influence of the hidden representations. We observe that smaller heterophilous graphs (e.g., Texas and Actor), optimal accuracy is achieved at low $\alpha=0.05$, implying that raw node features provide sufficient discriminative power. In contrast, on larger or homophilous graphs, better performance is observed when $\alpha=0.5$, reflecting the necessity of high-order hidden representations to capture more complex community structures. Moreover, for all datasets, the best results are gained at a larger $\alpha$, which demonstrates the effectiveness of excavating deep features.
Figure 4: Parameter sensitivity: Performance curves on five datasets with layers changing in $\{2,4,8,16,32\}$.
Figure 5 explores the effect of network depth $L$. For small graphs (Texas, Citeseer and Film), performance improves as the number of layers increases, since deeper networks are required to capture sufficient high-order information. In contrast, for large graphs (Tolokers and Flickr), the best performance is achieved with only two layers, indicating that shallow message passing already provides sufficiently discriminative representations. However, while APPNP can alleviate over-smoothing to some extent, it does not explicitly address this issue on these graphs; overcoming depth-related bottlenecks therefore remains an open direction for future research.
Figure 5: Parameter sensitivity: Performance curves on five datasets with layers changing in $\{2,4,8,16,32\}$.
Appendix D Broader Impact Statement